Suppose there are 2 routers in sequence between Host A and Host B, all of which use store-and-forward routing. What is the total end-to-end delay for a packet originating from Host A with destination Host B, under the following conditions. Each of the link transmission rates are 3.7 Mbps The total distance from Host A to Host B along its path of transmisison is 171.6 km The speed of propagation through the transmission medium is is 2.7 x 108 m/s The packet size is 3 KiB Remember that you must also uplink from Host A to the first router. Give answer in milliseconds, rounded to 1 decimal place, without units (e.g. for 0.12345 seconds you would enter "123.5" without the quotes).

How do you get to the answer of 20.6?
Answer #1

The path is A--R1--R2--B

Time taken to propagate through medium(The whole path) = (171.6 *1000 m )/(2.7*108 m/s)=0.63 ms

Transmission time per path between router = (3*1024*8 bits)/(3.7*106 bps) = 6.64 ms

Hence the total delay= 3 * Transmission Time taken between one path + Time taken to propagate through the medium = 6.64*3+0.63 = 20.55 ms

Do give a thumbs up